We are using the standard normal random variable is between 0 and Z table
I need to find the probability for
mean of 22 feet height and standard deviation is 4 feet
A. what is the probability that the height is greater than 17
B. height is less than 13 feet
C. the height is between 25 and 31 feet
I tried x-u/O to get z score and looked up in the table and all my answers have been wrong please help me understand how to find probability
Thank You
Beverly

Answer by stanbon(75887)

mean of 22 feet height and standard deviation is 4 feet
A. what is the probability that the height is greater than 17
z(17) = (17-22)/4 = -5/4
P(x > 17) = P(z > -5/4) = 0.8944
-----------------------------------
B. height is less than 13 feet
z(13) = (13-22)/4 = -9/4
P(x < 13) = P(z < -9/4) = 0.122
-------------------------------
C. the height is between 25 and 31 feet
---
Find the z value of 25 and of 31.
The find the probability that z is between those two values.
Ans: 0.2144
===================================
Cheers,
Stan H.
===================